Nivaisdead:

solve Quadraic equations by taking the square root 1). x2 =49 2).x2=73 3). x2+5=69 Help me please

BlkOps77:

1). x=7 and x=-7 2). ? 3). x=8 x=-8

Luigi0210:

@blkops77 wrote:
1). x=7 and x=-7 2). ? 3). x=8 x=-8
Care to elaborate on how you got those answers?

insatiable:

-Assuming x2 = x^2- 1) x^2 = 49 sqrt(x^2) = sqrt(49) x = 7 2) x^2 = 73 sqrt(x^2) = sqrt(73) The sqrt of 73 is an unfactorable number and is not a perfect square, so it must remain as sqrt(73) unless you are explicitly told to estimate/round so x = sqrt(73) 3) x^2 + 5 = 69 x^2 + 5 - 5 = 69 - 5 x^2 = 64 sqrt(x^2) = sqrt(64) x = 8 --The negative values of the afformentioned x solutions also apply, as a -# * a -# = a positive #

BlkOps77:

@luigi0210 wrote:
@blkops77 wrote:
1). x=7 and x=-7 2). ? 3). x=8 x=-8
Care to elaborate on how you got those answers?
I jus did this math in class

Extrinix:

@blkops77 wrote:
simple math atleast for me, idk wha math you guys do in school or wtv
As you were warned for, we try to teach users the content they post, not give them direct answers.